OpenStudy (anonymous):

obtain dy/dx: y=sec^3(x-x^2)^4

ganeshie8 (ganeshie8):

\[\large y = \sec^3\left(x-x^2\right)^4 = \large y = \left(\sec\left(x-x^2\right)^4\right)^3\]

ganeshie8 (ganeshie8):

use chain rule

ganeshie8 (ganeshie8):

\[ \large y' = 3\left(\sec\left(x-x^2\right)^4\right)^{3-1}\left(\sec\left(x-x^2\right)^4\right)'\]

ganeshie8 (ganeshie8):

keep unfolding using chain rule..